# Beyond I-Con: Exploring New Dimension of Distance Measures in Representation Learning

### v1

@@ Abstract
The Information Contrastive (I-Con) framework revealed that over 23 representation learning methods implicitly minimize KL divergence between data and learned distributions that encode similarities between data points. However, a KL-based loss may be misaligned with the true objective, and properties of KL divergence such as asymmetry and unboundedness may create optimization challenges. We present Beyond I-Con, a framework that enables systematic discovery of novel loss functions by exploring alternative statistical divergences and similarity kernels. Key findings: (1) on unsupervised clustering of DINO-ViT embeddings, we achieve state-of-the-art results by modifying the PMI algorithm to use total variation (TV) distance; (2) on supervised contrastive learning, we outperform the standard approach by using TV and a distance-based similarity kernel instead of KL and an angular kernel; (3) on dimensionality reduction, we achieve superior qualitative results and better performance on downstream tasks than SNE by replacing KL with a bounded f-divergence. Our results highlight the importance of considering divergence and similarity kernel choices in representation learning optimization.
Keywords Representation Learning  $\cdot$ Contrastive Learning  $\cdot$ Clustering  $\cdot$ Dimensionality Reduction
@@ 1 Introduction
The choice of optimization objective fundamentally determines the success of representation learning methods, yet the field has converged on a single statistical divergence measure without systematic exploration of alternatives. The Information Contrastive (I-Con) framework recently revealed that over 23 diverse representation learning methods all implicitly minimize KL divergence between data and learned distributions that encode similarities between data points alshammari2024icon .
This usage of KL divergence may lead to suboptimal optimization of representations. KL divergence is known to cause optimization issues due to properties such as asymmetry and the possibility of infinite values lazic2021 ; wassersteingan . Furthermore, since loss functions are only proxies for the actual goal of the task, optimizing for KL divergence may be misaligned with the true objective.
The natural question emerges: if representation learning methods can be unified under minimizing divergence of two distributions, what happens when we systematically explore alternative divergences and their interaction with similarity measures? We address this question through experimentation across multiple divergence-similarity combinations.
Contributions. We present Beyond I-Con, making the following contributions: (1) We generalize I-Con by replacing KL divergence with alternative f-divergences, revealing that KL is not unique in enabling meaningful feature optimization; (2) We systematically explore combinations of f-divergence and similarity kernel, uncovering novel loss functions with superior performance on unsupervised clustering, supervised contrastive learning, and dimensionality reduction. Specifically, we demonstrate that total variation distance when paired with distance similarity kernels achieves superior performance compared to KL-based approaches.
@@ 2 Background: I-Con Overview
The I-Con framework unifies representation learning methods by framing them as “minimizing the average KL divergence between two conditional ‘neighborhood distributions’ that define transition probabilities between data points,” which we index with $i=1,\ldots,N$ alshammari2024icon . $p(j\mid i)$ is typically a fixed “supervisory” distribution, and $q_{\phi}(j\mid i)$ are learnable transition probabilities typically calculated using similarities between features — see Figure 2a of the original paper for an illustration. The core I-Con loss function (Equation 1 of the original paper) is
@@ 3 Beyond I-Con Framework
3.1 Overview
We generalize the I-Con objective by replacing the KL divergence with any positive definite divergence $\mathcal{D}$:
@@ 3.1 Overview
We focus on $f$-divergences including KL, Total Variation (TV), Jensen-Shannon (JSD), and Hellinger because they are most directly comparable to KL as a measure of distance between distributions — some losses such as JSD directly remedy weaknesses of KL such as asymmetry.
3.2 Similarity Kernels
In the original I-Con paper, similarity kernels are defined on learned features which allows the transition probabilities of features to be defined ($q_{\phi}$). Two prominent kernels use squared euclidean distance and cosine similarity:
$k(z_{i},z_{j})=\frac{\exp(Cs(z_{i},z_{j}))}{\sum_{k\neq i}\exp(Cs(z_{i},z_{k}))}\quad\text{where }s(z_{i},z_{j})\in\{z_{i}\cdot z_{j},-\|z_{i}-z_{j}\|^{2}\}.$ (3)
In this paper, we create a novel loss for supervised contrastive learning khosla2020supervised by swapping the similarity kernel of the existing loss (cosine similarity) for a squared euclidean distance kernel. A similar procedure may be done on other losses such as SimCLR chen2020simple .
@@ 4.1 Unsupervised Clustering
We modify the Pointwise Mutual Information (PMI) clustering algorithm adaloglou2023exploringlimitsdeepimage by using different divergences than KL. We followed the same training setup as in the I-Con paper, which clusters DINO ViT embeddings caron2021emerging on ImageNet-1K deng2009imagenet by training a linear classifier for 30 epochs with a batch size of 4096, with a learning rate of $1\times 10^{-3}$ with an Adam optimizer. Data augmentation was applied to the training samples. Table 1 contains the results of this experiment. Clustering using PMI with TV outperforms the state-of-the-art on ViT-B/14 and ViT-L/14 embeddings.
@@
Contrastive Clustering li2021contrastive 47.35 55.64 59.84
SCAN van2020scan 49.20 55.60 60.15
TEMI niu2022temi 56.84 58.62
@@
Table 1: Comparison of methods on ImageNet-1K clustering with respect to Hungarian Accuracy. TV outperforms state of the art for ViT-B and ViT-L.
@@ 4.2 Supervised Contrastive Learning
We trained ResNet-50 he2016deep models on CIFAR-10 krizhevsky2009learning for 150 epochs, with a batch size of 4096 and learning rate of 5e-4. We systematically varied divergence measures and similarity kernels. We used the learned features to perform classification by either training an additional linear classifier on the features (linear probing) or applying $k$-nearest neighbors ($k$-NN). Results are shown in Table 2.
Divergence Angular-based Distance-based
KL 92.72 / 91.33 training collapse
TV 85.04 / 81.80 96.41 / 97.33
Hellinger 89.16 / 87.12 92.50 / 91.93
JSD 86.69 / 84.03 91.83 / 90.99
Across 5 seeds, KL divergence validation accuracy increases before collapsing to 10-20%.
Table 2: Linear probing / $k$-NN accuracy (%, top 7) on CIFAR-10. Bold indicates best performance.
The combination of TV and distance-based similarity kernel demonstrates greater performance than vanilla supervised contrastive learning (KL and angular kernel).
@@ 4.3 Dimensionality Reduction
We also ran SNE hinton2002sne using different divergences for CIFAR-10 to visualize qualitative differences between resulting image embeddings, as shown in Figure 1. Quantitative results from downstream classification are found in Table 3.
Divergence Linear Probe Test Accuracy $k$-NN Test Accuracy ($k=7$)
KL 57.63 74.00
TV 78.08 89.85
Hellinger 80.69 88.88
JSD 82.32 89.96
Table 3: Linear probing and $k$-NN accuracy (%) on CIFAR-10. Bold indicates best performance.
@@
5 Analysis and Discussion
Across the tasks of unsupervised clustering, supervised contrastive learning, and dimensionality reduction, we observe strong performance from TV in comparison to KL. We hypothesize this is because optimizing for KL overly penalizes placing dissimilar points farther apart in the feature space (KL diverges to infinity as $q(j\mid i)\to 0$). This causes different clusters or classes to crowd together and start overlapping with each other. For example, in dimensionality reduction, SNE hinton2002sne is known to have this crowding problem, which we also observe in Figure 1. Beyond I-Con provides a solution by replacing the KL divergence in the loss with other distances/divergences. TV, JSD, and Hellinger all penalize small values of $q(j\mid i)$ less than KL does, since they all remain bounded as $q(j\mid i)\to 0$.111 t-SNE van2008visualizing provides an alternative solution to the crowding problem by replacing $q(\cdot\mid i)$ with one where $q(j\mid i)$ does not decrease as quickly as the distance between $x_{i}$ and $x_{j}$ increases. Indeed, they all solve the crowding problem as seen in the resultant low-dimensional visualizations (Figure 1), and also significantly improve downstream task performance (Table 3).
We also hypothesize that KL-based losses may be prone to unstable gradients during training, similar to previous findings lazic2021 ; wassersteingan . In dimensionality reduction, our gradient norm plots during training (Figure 2) show large spikes in gradients near the beginning of training when the KL-based loss from vanilla SNE is used. In supervised contrastive learning (SupCon), a KL-based loss paired with a distance-based similarity kernel consistently results in training collapse.
Furthermore, our results in SupCon suggest that divergence cannot be chosen independently of the similarity measure because certain combinations lead to training instabilities, such as KL paired with a distance-based kernel. This may explain why existing methods use cosine similarity as previously all objectives were KL-based. We defer investigations on the exact interplay between divergence and similarity kernel to future work.
6 Limitations and Future Work
Our theoretical understanding of how divergence properties interact with similarity kernels to affect optimization landscapes remains incomplete. A rigorous analysis of gradient dynamics and loss surface geometry would strengthen the theoretical foundations of this work.
7 Conclusion
Beyond I-Con challenges representation learning methods’ implicit default usage of KL divergence by demonstrating that optimization with alternative statistical divergences can outperform KL-based methods and provides a proof-of-concept for novel loss discovery. Our exploration additionally suggests that the efficacy of novel losses also depends on instability and stability caused by specific loss-similarity kernel combinations. Finally, we demonstrate promise for this roadmap for novel loss discovery, as TV with distance-based similarity achieves state-of-the-art performance on unsupervised clustering on ImageNet-1K, beats vanilla supervised contrastive learning on CIFAR-10, and achieves superior results on CIFAR-10 with SNE.

### v2

# Beyond I-Con: A Roadmap for Representation Learning Loss Discovery

@@ Abstract
The Information Contrastive (I-Con) framework revealed that over 23 representation learning methods implicitly minimize KL divergence between data and learned distributions that encode similarities between data points. However, a KL-based loss may be misaligned with the true objective, and properties of KL divergence such as asymmetry and unboundedness may create optimization challenges. We present Beyond I-Con, a framework that enables systematic discovery of novel loss functions by exploring alternative statistical divergences. Key findings: (1) on unsupervised clustering of DINO-ViT embeddings, we achieve state-of-the-art results by modifying the PMI algorithm to use total variation (TV) distance; (2) supervised contrastive learning with Euclidean distance as the feature space metric is improved by replacing the standard loss function with Jenson-Shannon divergence (JSD); (3) on dimensionality reduction, we achieve superior qualitative results and better performance on downstream tasks than SNE by replacing KL with a bounded $f$-divergence. Our results highlight the importance of considering divergence choices in representation learning optimization.
keywords:
representation learning, contrastive learning, clustering, dimensionality reduction
editors: List of editors’ names
@@ 1 Introduction
The choice of optimization objective fundamentally determines the success of representation learning methods, yet the field has largely focused on a single statistical divergence measure without systematic exploration of alternatives. The Information Contrastive (I-Con) framework recently revealed that over 23 diverse representation learning methods all implicitly minimize KL divergence between data and learned distributions that encode similarities between data points (alshammari2024icon). The natural question emerges: if representation learning methods can be unified under minimizing the divergence between two distributions, what happens when we systematically explore alternative divergences?
Contributions. We present Beyond I-Con, making the following contributions: (1) We generalize I-Con by replacing KL divergence with alternative $f$-divergences, revealing that KL is not unique in enabling meaningful feature optimization; (2) We systematically explore $f$-divergences, uncovering novel loss functions with superior performance on unsupervised clustering, supervised contrastive learning, and dimensionality reduction.
@@ 2 Background: I-Con Overview
The I-Con framework unifies representation learning methods by framing them as “minimizing the average KL divergence between two conditional ‘neighborhood distributions’ that define transition probabilities between data points,” which we index with $i=1,\ldots,N$ (alshammari2024icon). $p(j\mid i)$ is typically a fixed “supervisory” distribution, and $q_{\phi}(j\mid i)$ are learnable transition probabilities typically calculated using similarities between features — see Figure 2a of the original paper for an illustration. The core I-Con loss function (Equation 1 of the original paper) is
@@ 3 Beyond I-Con Framework
We generalize the I-Con objective by replacing the KL divergence with any positive-definite divergence $\mathcal{D}$:
@@
We focus on $f$-divergences including KL, Total Variation (TV), Jensen-Shannon (JSD), and Hellinger because they are most directly comparable to KL as a measure of distance between distributions — some losses such as JSD directly remedy weaknesses of KL such as asymmetry and unboundedness.
@@ 4.1 Unsupervised Clustering
We modify the Pointwise Mutual Information (PMI) clustering algorithm (adaloglou2023exploringlimitsdeepimage) by using different divergences than KL. We followed the same training setup as in the I-Con paper, which clusters DINO ViT embeddings (caron2021emerging) on ImageNet-1K (deng2009imagenet) by training a linear classifier for 30 epochs with a batch size of 4096, a learning rate of $1\times 10^{-3}$, and the Adam optimizer (kingma2014adam). Data augmentation was applied to the training samples. Table 1 contains the results of this experiment. Clustering using PMI with TV outperforms the state-of-the-art on ViT-B/14 and ViT-L/14 embeddings.
@@
TEMI (adaloglou2023exploringlimitsdeepimage) 56.84 58.62
@@
Table 1: Comparison of methods on ImageNet-1K clustering with respect to Hungarian Accuracy. TV outperforms the state-of-the-art for ViT-B and ViT-L.
@@ 4.2 Supervised Contrastive Learning
Divergence Linear probe test acc. $k$-NN ($k=7$) test acc.
KL 90.03 $\pm$ 0.14 89.61 $\pm$ 0.13
TV 83.23 $\pm$ 0.18 82.95 $\pm$ 0.16
Hellinger 90.47 $\pm$ 0.08 90.40 $\pm$ 0.09
JSD 90.84 $\pm$ 0.11 90.62 $\pm$ 0.11
Table 2: Downstream classification accuracy from SupCon-learned features on CIFAR-10. Errors are standard errors of the mean over 5 seeds.
We trained ResNet-50 (he2016deep) models with supervised contrastive learning (khosla2020supervised) on CIFAR-10 (krizhevsky2009learning). We used a Euclidean distance metric on features and trained for 150 epochs with a batch size of 2048 and learning rate of 1e-3. We systematically varied divergence measures. We used the learned features to perform classification by training a linear probe or applying $k$-nearest neighbors ($k$-NN). Results are shown in Table 2.
Both Hellinger distance and Jenson-Shannon divergence (JSD) exhibit better performance than vanilla supervised contrastive learning (KL divergence). In particular, JSD achieves the best performance.
@@ 4.3 Dimensionality Reduction
We also ran SNE (hinton2002sne) with a CNN backbone on CIFAR-10 using different divergences to visualize qualitative differences between resulting image embeddings, as shown in Figure 1. Visually, while SNE with KL divergence creates highly overlapping clusters, the clusters resulting from SNE with the other divergences are more cleanly separated.
@@
5 Analysis and Discussion
Across the tasks of unsupervised clustering, supervised contrastive learning, and dimensionality reduction, we observe that a non-KL loss outperforms KL. We hypothesize this is because optimizing for KL overly penalizes placing dissimilar points farther apart in the feature space (KL diverges to infinity as $q(j\mid i)\to 0$). This causes different clusters or classes to crowd together and start overlapping with each other.
For example, in dimensionality reduction, SNE (hinton2002sne) is known to have this crowding problem, which we also observe in Figure 1. Beyond I-Con provides a solution by replacing the KL divergence in the loss with other distances/divergences. TV, JSD, and Hellinger all penalize small values of $q(j\mid i)$ less than KL does, since they all remain bounded as $q(j\mid i)\to 0$.111t-SNE (van2008visualizing) provides an alternative solution to the crowding problem by replacing $q(\cdot\mid i)$ with one where $q(j\mid i)$ does not decrease as quickly as the distance between $x_{i}$ and $x_{j}$ increases. Indeed, they all solve the crowding problem as seen in the resultant low-dimensional visualizations (Figure 1). Appendix B provides further analysis on the geometric arrangement of clusters in high- vs. low-dimensional spaces.
We also hypothesize that KL-based losses may be prone to unstable gradients during training, similar to previous findings (lazic2021; wassersteingan). For example, in dimensionality reduction, our gradient norm plots during training (Figure 2) show large spikes in gradients near the beginning of training when the KL-based loss from vanilla SNE is used.
6 Conclusion
In this work, we extended the I-Con representation learning framework with a new dimension — the type of $f$-divergence in the loss function. By experimenting with alternative divergences, we achieve state-of-the-art ImageNet-1K clustering, surpass vanilla supervised contrastive learning on CIFAR-10, and outperform SNE on CIFAR-10. Beyond I-Con challenges the default reliance on KL divergence in representation learning by showing that alternative statistical divergences can yield superior performance and serve as a basis for novel loss discovery.
Appendix A Supplementary Figures and Tables
@@
Appendix B Geometric Analysis of the Crowding Problem
In this appendix, we illustrate with an example why the KL divergence loss results in overcrowding in SNE.
Let the original high-dimensional space have $d$ dimensions. A $d$-dimensional space permits $d+1$ clusters/classes that are equidistant from each other — say they’re distance 1 from each other. Now, mapping these clusters to a reduced $d^{\prime}$-dimensional feature space ($d^{\prime}\ll d$) while ensuring the clusters are still separated by distance at least 1 means that some clusters will be very far from each other ($\Omega\big(d^{1/d^{\prime}}\big)$) (rogers1964packing). This incurs a high KL penalty as $q(j\mid i)\ll p(j\mid i)$ for data points $i$, $j$ from two clusters that are now much farther away each other. Thus, minimizing the KL loss inevitably leads to some clusters being brought too close together (crowding) in the low-dimensional feature space to prevent a high KL penalty from two clusters that are too far apart. Bounded divergences like TV, Hellinger, and JSD resolve this issue since $q(j\mid i)\ll p(j\mid i)$ incurs a lower penalty.